OpenStudy (anonymous):

Rewrite using scientific notation: 338. 3.38 × 10^1 3.38 × 10^2 3.38 × 10^3 3.38 × 10^4

OpenStudy (anonymous):

Let's simplify this number down a little bit 338 = 338 x 10^0 338 = 33.8 x 10^1 338 = 3.38 x 10^2 338 = 0.338 x 10^3 Does that make sense to you?

OpenStudy (anonymous):

yeah